A golfer puts her golf ball straight toward the hole. The ball's initial velocity is 3.97 m/s, and it accelerates at a rate of -0.6 m/s2. What is the maximum distance that the golfer can be from the hole in order to make it? Give your answer in meters.
Answer #1

using 3rd equation of motion

v^2 = u^2 + 2 a d

for maximum distance, v = 0 m/s

0 = 3.97^2 + 2* (-0.6)* d

d = 13.13 m

==========

do comment in case any doubt, will reply for sure Goodluck